From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Jan 17 2003 - 00:18:06 EST
------------------------------------------------------------ Date: Fri Jan 17 00:18:06 2003, Shift: 16:00-24:00 Supervisor: RP, with: JIJ Log book#13 pages 73 - 75 Shift Summary: ------------------------------------------------------------ 1. No beam during the shift. We were trying to find the "Vote To Start Store" web page, but we didn't manage. :) Pedestal and DVM runs. 2. Dana set the TOFW trigger 35 nsec wide. 3. Some scaler plots in supermon are empty (see attached pages in logbook). They looks different when comparing with run5326(Au+Au, Oct. 2001). Our assumption is that due to some hardware changes the empty channels don't go into the BRAHMS data stream. Watching the scalers during an attempt to ramp the blue beam there was very little background. The beam was beeing ramped and was lost a few seconds after we took the snapshots, so unfortunately we could not gauge the luminosity. See plots in logbook and elog. It may seem that the blue beem is rather clean. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Fri Jan 17 2003 - 00:23:02 EST